Sales Budget

Alger Inc. manufactures six models of leaf blowers and weed eaters. Alger's budgeting team is finalizing the sales budget for the coming year. Sales in units and dollars for last year follow:

Product Number Sold Price ($) Revenue
LB-1 14,700 32 $ 470,400
LB-2 18,000 20 360,000
WE-6 25,200 15 378,000
WE-7 16,200 10 162,000
WE-8 6,900 18 124,200
WE-9 4,000 22 88,000
Total $1,582,600

In looking over the previous year's sales figures, Alger's sales budgeting team recalled the following:

  1. Model LB-1 is a newer version of the leaf blower with a gasoline engine. The LB-1 is mounted on wheels instead of being carried. This model is designed for the commercial market and did better than expected in its first year. As a result, the number of units of Model LB-1 to be sold was forecast at 250% of the previous year's units.
  2. Models WE-8 and WE-9 were introduced on July 1 of last year. They are lighter versions of the traditional weed eater and are designed for smaller households or condo units. Alger estimates that demand for both models will continue at the previous year's rate.
  3. A competitor has announced plans to introduce an improved version of model WE-6, Alger's traditional weed eater. Alger believes that the model WE-6 price must be cut 40% to maintain unit sales at the previous year's level.
  4. It was assumed that unit sales of all other models would increase by 6%, prices remaining constant.

Required:

Prepare a sales budget by product and in total for Alger Inc. for the coming year. Do not include a multiplication symbol as part of your answer.

Alger Inc.
Sales Budget
For the Coming Year
Model Units Price Total Sales
LB-1 $ $
LB-2
WE-6
WE-7
WE-8
WE-9
Total $
Answer #1

Alger Inc.

Sales Budget

For the coming Year

Model Units Price

Total Sales

=Units×Price

LB-1

(Unit increase to 250%) [WN:1]

36,750 $32 $1,176,000

LB-2

(unit increase by 6%)[WN:2]

9,080

$20 $381,600

WE-6

(Price cut down by 40%)[WN:3]

25,200

$9

$226,800

WE-7

(Unit increased by 6%}[WN:4]

17,172

$10 $171,720
WE-8 6,900 $18 $124,200
WE-9 4,000 $22 $88,000
Total $2,168,320

WORKING NOTE[WN]:

1)No.of unit sold of LB-1= 14,700× 250%=36,750

2)No.of unit sales of LB-2= 18,000×(100+6)%=18,000×106%=19,080

3)Price of WE-6 is cut down by 40%= $15×(100-40)%=$15×60%=$9

4)Unit sales of WE-7= 16,200×(100+6)%=16,200(106)%=17,172